# Existence of Minimal Surfaces in Infinite Volume Hyperbolic $3$-manifolds

### v1

@@ 6.1. Construction of Barrier Surfaces in the Ends
The original statement of [Cos21, Lemma 2.3] is formulated for surfaces in the homotopy type of $\mathcal{T}$. However, by employing [Cos21, Lemma 2.2] in its proof, the result also extends to surfaces in the homology class of $\mathcal{T}$.
The main idea of this proof is that after fixing the compact set $\mathcal{K}$, with the area bound on $\Sigma$ coming from the fact that $\Sigma$ is homologically essential, one can bound the number of times the area minimizing surface $\Sigma$ can intersect the thick part $\mathcal{M}_{\geq\epsilon}$. Hence, the proof of this result applies to the following case.
Lemma 6.6.
Let $\mathcal{M}$ be a hyperbolic $3$-manifold with rank-1 cusps. Let $\mathcal{M}_{0}$ be the manifold obtained from $\mathcal{M}$ by removing rank-1 cusps. Let $\mathcal{E},\Sigma,\mathcal{K},\mathcal{Y}_{\mathcal{K}}$ be as in the previous lemma. Let $\widehat{\Sigma}$ be a connected component of $\mathcal{M}^{0}\cap\Sigma$ with $\widehat{\Sigma}\cap\mathcal{K}\neq\emptyset$. Then, there exists $C_{\mathcal{K}}>0$ such that $d_{\mathcal{K}}(\widehat{\Sigma})=\sup\{d(x,\mathcal{K})\mid x\in\widehat{\Sigma}\}<C_{\mathcal{K}}$.
Proof.
The main argument in 6.4 is to keep track of Margulis tubes near $\mathcal{K}$, and get an upper bound on the number of thick balls in $\Sigma$. As $\widehat{\Sigma}$ is connected, the same bound and argument apply here. ∎
By taking $\mathcal{K}=\alpha_{1}$ in the shrinkwrapping construction, and $d(\alpha_{1},\alpha_{2})>C_{\mathcal{K}}$, we can guarantee that shrinkwrapping surface $\Sigma$ can intersect only one of the geodesics. Unfortunately, the proof does not directly extend to the manifolds with rank-1 cusps, but by modifying the shrinkwrapping surface, we will obtain a barrier surface at each end.
To prove the existence of a barrier surface at each end, the last ingredient we need is a powerful result from geometric measure theory. Before we state the result, we recall that a compact surface $\Sigma$ with boundary in a Riemannian $3$-manifold $\mathcal{M}$ is called area minimizing surface if $\Sigma$ has the smallest area among all the surfaces in $\mathcal{M}$ with the same boundary. Similarly, a noncompact surface $\mathcal{S}$ is called area minimizing if any compact subsurface is an area minimizing surface.
Lemma 6.7.
(Existence of area minimizing surface in mean-convex domains [FF60, SSA77, Har77]) Let $\mathcal{M}$ be a homogeneously regular Riemannian manifold with strictly mean convex boundary. Let $\Gamma$ be a simple closed curve in $\partial\mathcal{M}$, and null-homologous in $\mathcal{M}$. Then there exists an area minimizing surface $\Sigma\subset M$ with $\partial\Sigma=\Gamma$ and $\Sigma$ is smooth away from its boundary and it is smooth around points of the boundary where $\Gamma$ is smooth.
Remark 6.8.
For the proof of 6.7, the existence result follows from [FF60], and the regularity results in [SSA77, Har77] extend from the ambient space being $\mathbf{R}^{n}$ to a mean convex manifold $\mathcal{M}$ as explained in [Whi91, Theorem 6.3]. The existence proof is very general and applies to any fixed homology class in $H_{2}(\mathcal{M},\partial\mathcal{M})$, see also  [Fed14].
Now, we are ready to prove our key result to construct finite area minimal surfaces.
Theorem 6.9.
[Existence of Barrier Surfaces] Let $\mathcal{M}$ be a hyperbolic $3$-manifold possibly with rank-$1$ cusps, and let $\mathcal{E}\cong\mathcal{T}\times[0,\infty)$ be an end of $\mathcal{M}$. Then, there exists a mean-convex or mean-concave barrier surface in $\mathcal{E}$.
Proof:  If $\mathcal{E}\simeq\mathcal{T}\times[1,\infty)$ is not cusped ($\mathcal{T}$ is a closed surface), then the proof follows from [Cos21, Theorem 3.1]. Therefore, we assume $\mathcal{M}$ contains rank-1 cusps and $\mathcal{E}$ is a cusped end. Let $\mathcal{M}^{0},\widehat{\mathcal{C}}_{\mathcal{M}},\mathcal{U}_{i},\widehat{\mathcal{U}},\mathcal{M}^{n},\widehat{\mathcal{U}}^{n}$ be as introduced in the beginning of the section.
Step 1. For a given pair of disjoint geodesics $\Lambda=\{\alpha_{1},\alpha_{2}\}$ in $\mathcal{E}$, there exists a complete shrinkwrapping surface $\Sigma_{\Lambda}$ of $\mathcal{E}$.
Proof of Step 1. Let $\mathcal{Y}_{\Lambda}$ be the special homology class as defined above, i.e., the space of complete surfaces in $\mathcal{M}$ in the homology class of $\widehat{\mathcal{T}}_{0}$ and separating $\alpha_{1}$ and $\alpha_{2}$ in $\mathcal{E}$. We show that there exists an area minimizing surface $\Sigma_{\Lambda}$ for $\mathcal{Y}_{\Lambda}$.
First, we will construct strictly-mean convex domains with two modifications on $\mathcal{M}^{n}$. The first is to drill out a small neighborhood of geodesics $\{\alpha_{1},\alpha_{2}\}$. The second is to change the metric near the boundary of the new manifold so that it will be strictly mean convex. Let $\{\epsilon_{n}\}$ be a monotone decreasing sequence with $\epsilon_{n}\to 0$. First, we drill out a small open neighborhood of the curves $\alpha_{1}$ and $\alpha_{2}$, let ~$\widetilde{\mathcal{M}}^{n}=\mathcal{M}^{n}-\mathcal{N}_{\epsilon_{n}}(\alpha_{1}\cup\alpha_{2})$. Next, we modify the metric in ~$\widetilde{\mathcal{M}}^{n}$ in $\mathcal{N}_{\epsilon_{n}}(\partial\widetilde{\mathcal{M}}^{n})$ so that the new manifold, say ^$\widehat{\mathcal{M}}^{n}$, is strictly mean convex. Notice that $\mathcal{N}_{\epsilon_{n}}(\partial\widetilde{\mathcal{M}}^{n})$ includes $\mathcal{N}_{\epsilon_{n}}(\partial\mathcal{M}^{n})$ and $\mathcal{N}_{2\epsilon_{n}}(\alpha_{i})-\mathcal{N}_{\epsilon_{n}}(\alpha_{i})$ for $i=1,2$.
Let $\mathcal{Y}_{\Lambda}$ be the special homology class defined at the beginning of the section, and let $\mathcal{F}\subset\mathcal{M}-\mathcal{N}_{\epsilon_{1}}(\alpha_{1}\cup\alpha_{2})$ be a complete surface in $\mathcal{Y}_{\Lambda}$. Let $\mathcal{F}^{n}=\mathcal{F}\cap\mathcal{M}^{n}$. For each $n$, define $\mathcal{Y}^{n}_{\Lambda}$ as a special homology class in $H_{2}(\widehat{\mathcal{M}}^{n},\partial\mathcal{M}^{n})$ analogously as follows. Note that $\partial\mathcal{M}^{n}\neq\partial\widehat{\mathcal{M}}^{n}$. Let $\mathcal{Y}^{n}_{\Lambda}$ be the space of compact surfaces $\mathcal{S}\subset\widehat{\mathcal{M}}^{n}$ with boundary $\partial\mathcal{S}\subset\partial\mathcal{M}^{n}$ such that $[\mathcal{S}]=[\mathcal{F}^{n}]\in H_{2}(\widehat{\mathcal{M}}^{n},\partial\mathcal{M}^{n})$.
Next, for each cusp $\mathcal{U}_{i}$, define a sequence of loops $\gamma_{i}^{n}=\gamma_{i}\times\{c_{0}\}\times\{n\}$ for some fixed sufficiently large $c_{0}>0$. Notice that the lengths of curves $|\gamma_{i}^{n}|\to 0$ as $n\to\infty$. Let $\Gamma^{n}=\bigcup_{i=1}^{m}\gamma_{i}^{n}$. Notice that $\Gamma^{n}\subset\partial\widehat{\mathcal{M}}^{n}$. By 6.7 and 6.8, there exists an area minimizing surface $\widehat{\Sigma}_{n}$ with $\partial\widehat{\Sigma}_{n}=\Gamma_{n}$ and in $\mathcal{Y}^{n}_{\Lambda}$ (separating $\alpha_{1}$ and $\alpha_{2}$). By 6.7, $\widehat{\Sigma}_{n}$ is a smooth surface in ^$\widehat{\mathcal{M}}^{n}$.
Let $\Sigma_{n}=\widehat{\Sigma}_{n}-\mathcal{N}_{\epsilon_{n}}(\partial\mathcal{M}^{n})$. Then, $\Sigma_{n}$ would be a compact surface with a boundary in $\mathcal{M}$. Notice that the surfaces $\{\Sigma_{n}\}$ may not be minimal in $\mathcal{M}$ as the original metric is modified in $\mathcal{N}_{2\epsilon_{n}}(\alpha_{i})-\mathcal{N}_{\epsilon_{n}}(\alpha_{i})$ for $i=1,2$.
Now, we apply the diagonal sequence argument to the sequence $\Psi_{0}=\{\Sigma_{n}\}$ to finish the proof of Step 1. Let $\Sigma^{j}_{n}=\Sigma_{n}\cap\mathcal{M}^{j}$. Then, for any $j$, $\{\Sigma^{j}_{n}\}$ is a sequence of smooth surfaces with boundary in $\mathcal{M}^{j}$. Let $\mathcal{R}^{j}_{n}=[\Sigma^{j}_{n}]$ be the corresponding integral current. For each $k>j$, $|\Sigma^{j}_{k}|<|\Sigma^{j}_{j}|+|\partial\mathcal{M}^{j}|$ where ||$\cdot$|| ⋅ | represent the area. This implies for $j$, $\{|\Sigma^{j}_{n}|\}$ is uniformly bounded. Hence, by Federer-Fleming compactness theorem for integral currents [FF60], we have a convergent subsequence of $\Psi_{1}=\{\Sigma^{1}_{\eta_{1}(i)}\}$ with $\lim_{i\to\infty}|\Sigma^{1}_{\eta_{1}(i)}|=\inf\{|\Sigma^{1}_{n}|\}$ where $\eta_{1}:\mathcal{N}\to\mathcal{N}$ a monotone increasing function. Next, for each $j>1$, we iteratively apply the Federer-Fleming compactness theorem and obtain a subsequence $\Psi_{j}=\{\Sigma_{\eta_{j}(i)}\}$ of $\Psi_{j-1}$ with $\eta_{j}:\mathbb{N}\to\mathbb{N}$ such that $\lim_{i\to\infty}|\Sigma^{j}_{\eta_{j}(i)}|=\inf\{|\Sigma^{j}_{n}|\}$. Let $\mathcal{R}^{j}=\lim_{i\to\infty}\mathcal{R}^{j}_{\eta_{j}(i)}$ and let $\Sigma^{j}$ be the support of $\mathcal{R}^{j}$. By regularity theorem [Fed14], for each $j$, $\Sigma^{j}$ is a smooth surface with boundary in $\mathcal{M}^{j}$. Furthermore, as $\epsilon_{n}\to 0$, $\mathcal{N}_{\epsilon_{n}}(\alpha_{i})$ shrinks to $\alpha_{i}$, and hence for any $j$, as $\Sigma^{j}_{n}$ is an area minimizing surface in $\mathcal{M}^{j}\cap\widehat{\mathcal{M}}^{n}$ where the metric is modified only in $\mathcal{N}_{2\epsilon_{n}}(\alpha_{i})-\mathcal{N}_{\epsilon_{n}}(\alpha_{i})$ in $\mathcal{M}^{j}$, the limit $\Sigma^{j}$ is indeed area minimizing surface in $\mathcal{Y}^{j}_{\Lambda}$. By [CG06, Lemma 1.31], $\Sigma^{j}$ is a piecewise smooth minimal surface except the coincidence set $\Sigma^{j}\cap\alpha_{i}$, and $\Sigma^{j}$ has tangent planes from each side along the coincidence set.
Notice that $\Sigma^{j}\subset\Sigma^{j+1}$ by construction. Let $\Sigma=\bigcup_{j}\Sigma^{j}$. As $\partial\Sigma_{n}$ escapes to infinity, $\Sigma$ is a complete minimal surface in $\mathcal{M}$. By construction, $\Sigma$ is separating $\alpha_{1}$ and $\alpha_{2}$, and in the same homology class of $\widehat{\mathcal{T}}_{1}$. By construction, $\Sigma$ is an area minimizing surface in $\mathcal{Y}_{\Lambda}$. The proof of Step 1 follows.
Step 2: $\mathcal{E}$ contains a mean-convex or mean-concave barrier surface.
Proof of Step 2. In Step 1, we constructed a complete shrinkwrapping surface in the homology class of $\mathcal{T}$ for a given geodesic pair $\{\alpha_{1},\alpha_{2}\}$ in an end $\mathcal{E}\simeq\mathcal{T}\times[1,\infty)$. However, this surface may intersect both geodesics, which prevents us from classifying $\Sigma$ as either a convex or concave shrinkwrapping surface. In this step, we will use 6.4 to demonstrate that if $\alpha_{1}$ and $\alpha_{2}$ are sufficiently far from each other, then we will have a barrier surface. Note that at this step, if there is a lower bound for the injectivity radius of the manifold ($\rho_{\mathcal{M}}>\epsilon_{0}$), one can employ much stronger results [Cos21, Lemma 2.1, 2.2] which imply uniform diameter bound on the minimal surfaces ($\mbox{diam}(S)<C_{\epsilon_{0}}$) without any reference to the compact set $\mathcal{K}$.
Figure 2. Mean-Convex Barrier Surface via Surgery. In the figure above, $\Sigma$ (Green) and $\widehat{\mathcal{T}}_{1}$ (Purple) are homologous with $\partial\mathcal{X}=\Sigma\cup\widehat{\mathcal{T}}_{1}$. If $\Sigma\cap\mathcal{M}^{0}$ is not connected (say $\mathcal{S}_{1}\cup\mathcal{S}_{2}$), the bounded diameter lemma does not apply to the away component $\mathcal{S}_{2}$ and hence $\mathcal{S}_{2}\cap\alpha_{2}$ may be nontrivial. To avoid this, we remove $\mathcal{S}_{2}$ and replace with $\mathcal{Z}_{2}$ in $\Sigma$ yielding a mean convex barrier surface $\widetilde{\Sigma}=(\Sigma-\mathcal{S}_{2})\cup\mathcal{Z}_{2}$, where $\widetilde{\Sigma}\cap\alpha_{2}=\emptyset$.
For the compact set $\mathcal{K}_{0}=\overline{\mathcal{T}}\times\{1\}$, let $C_{0}$ be the constant given by 6.4 (or 6.6 if rank-1 cusp exists). Consider a geodesic $\alpha_{1}$ in $\mathcal{E}$ that is sufficiently far from $\mathcal{K}_{0}$, specifically $d(\alpha_{1},\mathcal{K}_{0})=\inf\{d(x,y)\mid x\in\alpha_{1},y\in\mathcal{K}_{0}\}>C_{0}$. Again, for the compact set $\mathcal{K}_{1}=\alpha_{1}$, let $C_{1}$ be the constant given by 6.4, assuming $C_{1}>C_{0}$. Let $\alpha_{2}$ be a geodesic in $\mathcal{M}^{0}$ with $d(\alpha_{2},K_{0})>C_{0}+C_{1}$. Let $\Sigma$ be the shrinkwrapping surface for $\{\alpha_{1},\alpha_{2}\}$ constructed in Step 1. If $\Sigma$ does not intersect $\alpha_{1}$, the proof is complete. If $\Sigma\cap\alpha_{1}\neq\emptyset$, We claim that the shrinkwrapping surface $\Sigma$ can be modified into a mean convex barrier surface in $\mathcal{E}$.
Now, consider the intersection $\Sigma\cap\mathcal{M}^{0}=\Sigma^{0}$. If $\Sigma^{0}$ is connected, by 6.4 (or 6.6), $d_{\alpha_{1}}(\Sigma^{0})=\sup\{d(x,\alpha_{1})\mid x\in\Sigma^{0}\}<C_{1}$. Since $\alpha_{2}\subset\mathcal{M}^{0}$, this implies $\Sigma\cap\alpha_{2}=\Sigma^{0}\cap\alpha_{2}=\emptyset$, and the proof follows. Let $\mathcal{S}_{1}$ be the connected component of $\Sigma^{0}$ with $\mathcal{S}_{1}\cap\alpha_{1}\neq\emptyset$. Let $\Sigma^{0}-\mathcal{S}_{1}=\mathcal{S}_{2}$ where $\mathcal{S}_{2}$ is possibly disconnected (See Figure 2).
Let $\widehat{\mathcal{T}}_{1}$ be the complete cusped surface in the end $\mathcal{E}$ as described earlier. By construction, $\Sigma$ and $\widehat{\mathcal{T}}_{1}$ are homologous in $\mathcal{E}$. Let $\mathcal{X}\subset\mathcal{E}$ be codimension-00 domain in $\mathcal{M}$ with $\partial\mathcal{X}=\Sigma\cup\widehat{\mathcal{T}}_{1}$. Let $\mathcal{X}^{0}=\mathcal{X}\cap\overline{\mathcal{M}^{0}}$. Hence, $\partial\mathcal{X}_{0}=(\overline{\mathcal{T}\times{1}})\cup\overline{\mathcal{S}_{1}}\cup\overline{\mathcal{S}_{2}}\cup\mathcal{Z}$ where $\mathcal{Z}$ is a compact subsurface of $\partial\widehat{\mathcal{U}}$, the union of infinite cylinders in the boundary of rank-$1$ cusps. By assumption, $\mathcal{S}_{1}$ is separating in $\mathcal{M}^{0}$, therefore $(\overline{\mathcal{T}\times{1}})\cup\overline{\mathcal{S}_{1}}$ must separate a component of $\mathcal{X}^{0}$, say $\mathcal{X}^{0}_{1}$. In other words, $\partial\mathcal{X}^{0}_{1}=(\overline{\mathcal{T}\times{1}})\cup\overline{\mathcal{S}_{1}}\cup\mathcal{Z}_{1}$ where $\mathcal{Z}_{1}\subset\mathcal{Z}$ is the compact surface with $\mathcal{Z}_{1}=\mathcal{Z}\cap\partial\mathcal{X}^{0}_{1}$. Let $\mathcal{Z}_{2}=\mathcal{Z}-\mathcal{Z}_{1}$ and $\mathcal{X}^{0}-\mathcal{X}^{0}_{1}=\mathcal{X}^{0}_{2}$. Then, $\partial\mathcal{X}^{0}_{2}=\overline{\mathcal{S}_{2}}\cup\mathcal{Z}_{2}$. Note that all $\mathcal{X}^{0}_{2},\mathcal{S}_{2}$ and $\mathcal{Z}_{2}$ might be disconnected.
Let $\widetilde{\Sigma}=(\Sigma-\mathcal{S}_{2})\cup\mathcal{Z}_{2}$. Recall that $\widehat{\mathcal{U}}=\bigcup_{i=1}^{m}\mathcal{U}_{i}$, and each rank-$1$ cusp $\mathcal{U}_{i}$ is a convex domain in $\mathcal{M}$. Notice that $\widetilde{\Sigma}$ is still homologous to $\Sigma$ as $\Sigma-\widetilde{\Sigma}$ bounds $\mathcal{X}^{0}_{2}$. Then, let $\mathcal{E}-\widetilde{\Sigma}$ be two components $\mathcal{E}_{in}$ and $\mathcal{E}_{out}$ where $\mathcal{E}_{in}$ has finite volume. Furthermore, $\widetilde{\Sigma}\cap\mathcal{M}^{0}=\mathcal{S}_{1}\subset\partial\overline{\mathcal{E}_{in}}$. This implies the mean curvature vector along $\mathcal{Z}_{2}\subset\partial\widehat{\mathcal{U}}$ points toward $\mathcal{E}_{in}$ and nonsmooth parts along $\partial\mathcal{Z}_{2}$ have dihedral angles less than $\pi$ as $\widetilde{\Sigma}$ stays in the same side of $\partial{\mathcal{M}^{0}}$ near $\mathcal{Z}_{i}$. This shows that $\widetilde{\Sigma}$ is a mean-convex barrier surface in $\mathcal{E}$. The proof follows.
6.2. Existence of Finite Area Minimal Surfaces
In this part, we prove the existence of a finite area minimal surface for manifolds with rank-1 cusps, when one of the ends contains a convex barrier surface.
Theorem 6.10.
Let $\mathcal{M}$ be an infinite volume complete hyperbolic $3$-manifold with rank-1 cusps. Assume $\mathcal{M}$ is topologically not a handlebody, nor a product manifold ($\mathcal{M}=\mathcal{S}\times\mathbf{R}$ for a closed surface $\mathcal{S}$). If one of the ends contains a convex barrier surface, then $\mathcal{M}$ contains a finite area, complete, embedded, stable minimal surface $\mathcal{S}$.
Proof:  The main idea is to use barrier surfaces in each end to obtain finite area complete minimal surfaces. If at least one of the ends contains a convex barrier surface, then we will create a minimal surface in the homology class of that barrier surface.
Let $\mathcal{M}$ be an infinite volume complete hyperbolic $3$-manifold with rank-1 cusps, and $\mathcal{M}^{0}$ be the manifold obtained from $\mathcal{M}$ by removing all rank-$1$ and rank-$2$ cusps. Let $\widehat{\mathcal{C}}_{\mathcal{M}}$ be the corresponding relative compact core of $\mathcal{M}$ (see Section 2.3). Assume $\mathcal{M}$ has $N$-ends, $\{\mathcal{E}_{1},\mathcal{E}_{2},\dots,\mathcal{E}_{N}\}$ where $\mathcal{E}_{i}\simeq\mathcal{T}_{i}\times[0,\infty)$ for a closed (if $\mathcal{E}_{i}$ is non-cusped) or an open surface (if $\mathcal{E}_{i}$ is cusped) $T_{i}\subset\partial\widehat{\mathcal{C}}_{\mathcal{M}}$. If an end $\mathcal{E}_{i}$ is geometrically finite, we simply take the corresponding boundary component $\mathcal{S}_{i}$ in the boundary of the convex core as the convex barrier surface. If not, let $\mathcal{S}_{i}$ be the corresponding barrier surface in $\mathcal{E}_{i}$ given by 6.9.
For simplicity, let $\mathcal{S}_{1}$ be a convex barrier surface in the end $\mathcal{E}_{1}$, where $\mathcal{E}_{1}\simeq\mathcal{T}_{1}\times[0,\infty)$. If $\mathcal{E}_{1}$ is a non-cusped end, the result follows directly from [Cos21, Theorem 5.9 - Step 3] with slight modification as follows. In this case, the only difference would be some of the remaining ends might have rank-$1$ cusp. However, since we assume $\mathcal{M}$ is not a topological handlebody, the area minimizing surface in the homology class of $\mathcal{S}_{1}$ would be non-degenerate (Area $>0$). Furthermore, since $\mathcal{M}$ is not a product manifold, it would have a nontrivial intersection with the (relative) compact core by 6.4 (or 6.6). This shows we have a finite area minimal surface $\Sigma$ intersecting the compact core of $\mathcal{M}$, or a closed minimal surface completely in $\mathcal{E}_{1}$. Hence, having rank-$1$ cusps in the other ends does not affect the proof. Therefore, in the remaining part of the proof, we assume that the end $\mathcal{E}_{1}$ is cusped.
Using $\mathcal{S}_{1}$, we will construct a finite volume mean convex domain $\Omega$ and define a sequence of compact area minimizing surfaces $\Sigma_{n}\subset\Omega$ in the homology class of $\mathcal{S}_{1}$, concluding the proof by taking the limit of this sequence.
Let $K>0$ be the constant provided by the Bounded Diameter Lemma (6.4) for the relative compact core $\widehat{\mathcal{C}}_{\mathcal{M}}$. This ensures that any minimal surface $\Sigma$ intersecting $\widehat{\mathcal{C}}_{\mathcal{M}}$ must satisfy $\Sigma\cap\mathcal{M}^{0}\subset\mathcal{N}_{K}(\widehat{\mathcal{C}}_{\mathcal{M}})$. To construct a finite-volume, mean convex domain $\Omega$, we will cut sufficiently deep into each of the remaining ends $\{\mathcal{E}_{2},\mathcal{E}_{3},\dots,\mathcal{E}_{N}\}$, and modify the metric near the boundary to ensure mean convexity. Let ^$\widehat{\mathcal{E}}_{1}$ denote the finite volume component of $\mathcal{E}_{1}-\mathcal{S}_{1}$. For any $i\geq 2$, define ^$\widehat{\mathcal{E}}_{i}=\mathcal{T}_{i}\times[0,\mathcal{C}_{i})$ as a finite volume domain, with $d(\mathcal{T}_{i}\times\{\mathcal{C}_{i}\},\widehat{\mathcal{C}}_{\mathcal{M}})>2K$. Let $\Omega=(\bigcup_{i=1}^{N}\widehat{\mathcal{E}}_{i})\cup\widehat{\mathcal{C}}_{\mathcal{M}}\cup_{j}\mathcal{V}_{j}$, where $\{\mathcal{V}_{j}\}$ are all rank-2 cusps in $\mathcal{M}$. By modifying the metric near $\partial\Omega-\mathcal{S}_{1}$, we construct a complete, finite volume, mean convex domain $\widehat{\Omega}$. Notice that $\widehat{\Omega}$ is homeomorphic to $\Omega$, but its metric is altered only in the region near $\partial\Omega-\mathcal{S}_{1}$.
Recall that $\widehat{\Omega}$ has finite volume. By following the notation at the beginning of the section, let $\{\mathcal{U}_{i}\}_{i=1}^{m}$ be all the rank-1 cusps in $\mathcal{M}$ with a parametrization $\mathcal{U}_{i}\simeq\gamma_{i}\times(-\infty,\infty)\times[0,\infty)$ where $\gamma_{i}$ is a core curve for $\mathcal{U}_{i}$. Let $\mathcal{U}_{i}^{n}=\gamma_{i}\times(-\infty,\infty)\times(n,\infty)$. Let $\widehat{\mathcal{U}}^{n}=\bigcup_{i=1}^{m}\mathcal{U}_{i}^{n}$ union of subsets of $\bigcup_{i=1}^{m}\mathcal{U}_{i}$ escaping to infinity. Let $\{\widehat{\Omega}^{n}\}=\widehat{\Omega}-\widehat{\mathcal{U}}^{n}$, which induce a compact exhaustion of $\widehat{\Omega}$, i.e., $\widehat{\Omega}^{n}\subset\widehat{\Omega}^{n+1}$ and $\widehat{\Omega}=\bigcup_{n=1}^{\infty}\widehat{\Omega}^{n}$. By modifying the metric in $\epsilon$-neighborhood of $\mathcal{N}_{\epsilon}(\partial\mathcal{U}^{n})\cap\widehat{\Omega}^{n}$ for a small $\epsilon>0$, we obtain a sequence of compact mean convex domains $\widetilde{\Omega}^{n}$. Now, define $\Gamma^{n}=\mathcal{S}_{1}\cap\partial\mathcal{U}^{n}$. Let $\Sigma_{n}$ be a compact area minimizing surface with boundary $\Gamma^{n}$ in $\widetilde{\Omega}^{n}$, whose existence is guaranteed by 6.7, as the manifold $\widetilde{\Omega}^{n}$ is mean-convex.
We claim that $\widehat{\Sigma}_{n}=\Sigma_{n}-\mathcal{N}_{\epsilon}(\partial\Sigma_{n})$ is a compact area minimizing surface with boundary in $\mathcal{M}$ with the original metric, i.e., $\Sigma_{n}$ can not wander away to interact with remaining boundary components of $\widetilde{\Omega}^{n}$. Recall that $\mathcal{M}$ is not a handlebody or geometrically infinite product manifold. First, since $\mathcal{M}$ is not a handlebody, the area minimizing surface in the homology class of $\mathcal{S}_{1}$ would be non-degenerate (Area $>0$). Next, again $\mathcal{M}$ is not a topological handlebody, it cannot be topologically equivalent to a product manifold ×$\mathcal{F}\times\mathbf{R}$ for a cusped surface $\mathcal{F}$. This implies $\mathcal{M}$ has at least three ends (cusped or non-cusped), and any surface in the homology class of $\Sigma$ cannot lie entirely within one of the other ends $\{\mathcal{E}_{2},\mathcal{E}_{3},\dots,\mathcal{E}_{n}\}$. In other words, $\Sigma_{n}$ is either fully contained in $\mathcal{E}_{1}$ or it has a nontrivial intersection with the relative compact core $\widehat{\mathcal{C}}_{\mathcal{M}}$. If $\Sigma_{n}\cap\widehat{\mathcal{C}}_{\mathcal{M}}\neq\emptyset$, then by 6.4, $\Sigma_{n}\cap\mathcal{M}^{0}\subset\mathcal{N}_{K}(\widehat{\mathcal{C}}_{\mathcal{M}})$. This implies that $\Sigma_{n}$ is away from the regions in $\widehat{\Omega}$ where the metric has been modified to ensure mean convexity. Therefore, $\widehat{\Sigma}_{n}$ is indeed a compact area minimizing surface with boundary in $\mathcal{M}$ with the original metric.
[…]

### v2

@@ 6. Hyperbolic $3$-Manifolds with Rank One Cusps
In the case of quasi-Fuchsian hyperbolic manifolds homeomorphic to the product of a cusped surface with $\mathbb{R}$, the paper [HLS23] proved existence of complete embedded finite area minimal surface. Such manifolds satisfy the assumptions of the first sentence of the theorem but not the second.
Proof:  Let $\Sigma$ be the boundary of the convex core of $\mathcal{M}$ in $\mathcal{E}$. The main idea is to find an area minimizer starting with $\Sigma$ and using $\Sigma$ as a barrier, where the topological assumption are to prevent minimizing sequences from escaping to infinity. As the results of [MSY82] require a closed surface, we will modify and double the metric of $\mathcal{M}$ to obtain a sequence of manifolds ^$\widehat{\mathcal{M}}_{n}$ with area minimizers $\mathcal{S}_{n}$ and so produce the desired minimal surface $\mathcal{S}$ as a limit of $\mathcal{S}_{n}\in\widehat{\mathcal{M}}_{n}$.
Let $\{\mathcal{U}_{i}\}_{i=1}^{m}$ be the rank-1 cusp neighborhoods for all of the cusp ends of $\Sigma$. Without loss of generality we take $\{\mathcal{U}_{i}\}_{i=1}^{m}$ so that $\{\Sigma\cap\mathcal{U}_{i}\}_{i=1}^{m}$ are all totally geodesic components (see [CEG06, Chapter 4]). We claim that the pleating locus on $\Sigma$ is disjoint from $\cup_{i=1}^{m}\mathcal{U}_{i}$. Indeed, as $\Sigma$ is a finite area pleated surface we can take cusp neighborhoods so that every $2$ dimensional totally geodesic strata close to the cusp accumulates into it. In the universal cover, each such strata lifts to a totally geodesic region $R$ with totally geodesic boundary so that the boundary at infinity $\partial_{\infty}R$ contains a rank-1 parabolic fix point $p_{0}$ of $\Gamma$. As $R$ is given by a supporting plane of the convex hull of $\Gamma$ for any of its points, it must be invariant by the rank-1 parabolic subgroup of $\Gamma$ that fixes $p_{0}$. Moreover any such region $R$ must belong to the same invariant totally geodesic plane, from which the claim follows.
For each $\mathcal{U}_{i}$ take a parametrization by horocylinders $\mathbb{S}^{1}\times\mathbb{R}\times\mathbb{R}^{+}$ (where every slice $S^{1}\times\mathbb{R}\times\{t\}$ is the quotient of an horoball by the corresponding rank-1 parabolic subgroup of $\Gamma$). Let $\mathcal{U}^{n}_{i}\subset\mathcal{U}_{i}$ given by $\mathbb{S}^{1}\times\mathbb{R}\times]n,+\infty[$ and let $\mathcal{M}_{n}:=\mathcal{M}\setminus\left(\cup_{i=1}^{n}\mathcal{U}^{n}_{i}\right)$. Let $g$ be the hyperbolic metric in $\mathcal{M}$ and consider a conformal change of metric $\tilde{g}_{n}=e^{2u_{n}}g$ in $\mathcal{M}_{n}$ with $u_{n}$ satisfying the following properties
  1. (1)
    $u_{n}$ is a smooth function supported in $\mathbb{S}^{1}\times\mathbb{R}\times[n-1,n+1]$.
  2. (2)
    In $\mathbb{S}^{1}\times\mathbb{R}\times[n-1,n]$ the level sets of $u_{n}$ are given by horocylinders $S^{1}\times\mathbb{R}\times\{t\}$.
  3. (3)
    $\partial_{t}u_{n}\leq 1$ for $t\in[n-1,n]$, with equality at $\mathbb{S}^{1}\times\mathbb{R}\times\{n\}$.
It is straightforward to verify that such conformal changes exist.
The mean curvature of a surface under a conformal change of metric is given by
$H_{\tilde{g}}=e^{-u}\left(H_{g}-D_{n}u\right),$
where $n$ is the normal of a surface at a given point. Since by construction we have that $\Sigma$ is orthogonal to the level sets of $u_{n}$, it follows that $\Sigma$ is mean convex with respect to $g_{n}$.
By the third condition on $u_{n}$ we have that $\{S^{1}\times\mathbb{R}\times\{t\}\}_{t\in[n-1,n]}$ is a mean concave foliation with $S^{1}\times\mathbb{R}\times\{n\}$ minimal (i.e., the mean curvature vectors of the $\{S^{1}\times\mathbb{R}\times\{t\}\}_{t\in[n-1,n)}$ point towards $S^{1}\times\mathbb{R}\times\{n\}$.) Moreover, as one can check that under our assumptions each horocylinder stays umbilic, it follows that $S^{1}\times\mathbb{R}\times\{n\}$ is in fact totally geodesic in $\tilde{g}_{n}$.
[…]